OpenStudy (anonymous):

I need help with this word problem...Please help. :) Solve the problem. Be sure to show all of your work. When a ball is thrown upward at a speed of 19 m/s, its height s above the ground (in meters) after t seconds is given by the formula s = 19t - 4.9t^2 Find the height of the ball after 3 seconds.

OpenStudy (anonymous):

s= your position, aka your height...so just plug in 3 for t. s=19(3)-4.9(9)=57-44.1=12.9meters